Kofi is five times as old as kwaku. In ten years time kofi will be three times as old as kwaku. How old are the boys.

Answer by ikleyn(52776)

.
Kofi is five times as old as kwaku. In ten years time kofi will be three times as old as kwaku.
How old are the boys.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Let Kwaku age be x years.

Then Kofi age is 5x years.


In ten years Kofi's age will be  5x+10  years;  Kwaku age will be (x+10) years.

From the condition, we have THIS equation


    5x + 10 = 3*(x+10)


Simplify and find x

    5x + 10 = 3x + 30

    5x - 3x = 30 - 10

      2x    =    20

       x    =    20/2 = 10.


Thus Kwaku's age is 10 years.   Kofi's age is 5*10 = 50 years.      ANSWER

Solved and explained.

----------------

It is hardly possible to consider 50-years old Kofi as a boy . . .